# A continuous boostlet transform for acoustic waves in space–time

### v1

@@ 5 Uncertainty principles and minimizers
Following similar steps to Dahlke et al. [65], we can try to find polynomial solutions by setting $b:=-i/(2\eta)$, which reduces the above differential equation to
[…]
$h(\rho)=h_{0}+h_{2}\rho^{2},$
are solutions to (40) provided the coefficients $h_{0},h_{2}\in\mathbb{C}$, satisfy:
$04h_{2}+\tilde{q}h_{0}=4h_{2}-\frac{2i}{\eta}h_{0}=$ 0 .
Lastly, the Mexican hat wavelet is recovered by setting
$h_{0}=2,\quad h_{2}=-2b=\frac{i}{\eta},$
which gives
$\psi(x,t)=(2-2b\rho^{2})e^{-b\rho^{2}}.$
This result is analogous for the ansatz $\psi(x,t)=g(\rho)$ on the space-like cone, i.e., $\rho=\sqrt{x^{2}-t^{2}},x^{2}-t^{2}>0$, by setting $\hat{T}$ as minus the d’Alembert operator in 2D space–time, $\hat{T}(\psi)=-\Box\psi$. ∎
Remark 5.1.
The uncertainty minimizer (36) is square-integrable and admissible. This follows immediately from the definition of the Mexican hat wavelet as a function of the space–time cone and its associated Fourier spectrum. In other words, the minimizer is supported away from the radiation cone $\omega^{2}=k^{2}$.
Remark 5.2.
The infinitesimal generators of the boostlet group can be interpreted as a canonical basis for the Lie algebra. The uncertainty minimizers of Theorem 4 partially preserve weak invariance with respect to a change of the basis. This remark is akin to the uncertainty minimizers of 2D wavelets [65]. For example, let us consider the nontrivial linear combinations
$T_{1}:=d_{1}T_{\theta}+d_{2}T_{\tau_{x}},\quad S_{1}:=e_{1}T_{\theta}+e_{2}T_{\tau_{t}}.$
Using a similar calculation to that in the proof of Theorem 4, it follows that
$g^{\prime\prime}(\rho)+\left(\frac{1}{\rho}-\frac{i\rho}{\hat{\eta}}\right)g^{\prime}(\rho)+\hat{q}g(\rho)=0,$
where
[…]
and, consequently, $g(\rho)=(2-2\hat{b}\rho^{2})e^{-\hat{b}\rho^{2}}$, with $\hat{b}=-i/(2\hat{\eta})$.
6 Conclusion
This paper introduces a representation system for spatiotemporal acoustic waves that encodes the dispersion relation of the propagation medium using the Poincaré group and isotropic dilations. Several interesting results concerning admissibility conditions and uncertainty principles are discussed and interpreted physically. A sparsity analysis with experimentally measured acoustic fields shows that boostlet expansions exhibit a significantly faster coefficient decay compared to various benchmarks, including wavelets, curvelets, and wave atoms. It is noteworthy that boostlets are optimally concentrated with respect to dilation and boosting, since the infinitesimal generators of the dilation and boost operators commute, i.e., $\mathrm{var}_{a}\mathrm{var}_{\theta}\geq 0$. This expression suggests the existence of a continuum of wave shapes where $\mathrm{var}_{a}\to 0$ or $\infty$, and $\mathrm{var}_{\theta}\to 0$ or $\infty$, which appears to include canonical wave solutions such as plane waves and spherical waves. The most straightforward case to show heuristically now is when both $\mathrm{var}_{a}\to 0$ and $\mathrm{var}_{\theta}\to 0$, which can be recognized as a Dirac delta in wavenumber–frequency space, i.e., a stationary planar wave. Future works concern formalizing the continuum of wave shapes, deriving nonlinear approximation rates of boostlets, the operator for which they are eigenfunctions, extensions to higher spatial dimensions, and a digitally faithful implementation of the transform. In a discrete transform implementation, the isotropic dilations could be chosen such that the wavenumber–frequency bands follow hyperbolic frequency octave bands, thus having physical grounds from psychoacoustics. We deem the connections between sparse representations of acoustic fields, special relativity, quantum mechanics, and harmonic analysis fascinating. Pursuing the “momentum” of geometric multi-scale analysis [78], we hope this study stimulates encoding further physics into sparse representation systems.
Acknowledgments
E.Z. is financially supported by the Swedish Research Council (Vetenskapsrådet) under Grant Agreement No. 2020-04668. M.L. is partly supported by the Swedish Research Council (Vetenskapsrådet) under Grant Agreement No. 2022-03032. The authors would like to thank D. Labate, F. Lizzi, U.P. Svensson, E. Fernandez-Grande, O. Öktem, and C.E. Yarman for their valuable suggestions and discussions.

### v2

@@ Acknowledgments
$2^{2\ell}e^{-2|n|\delta}\sum_{m}\left|\int_{[0,2^{-\ell}e^{|n|\delta}]^{2}}e^{2\pi i2^{\ell}e^{-|n|\delta}m^{\text{T}}\xi}g(\xi)\mathrm{d}\xi\right|^{2}=\int_{[0,2^{-\ell}e^{|n|\delta}]^{2}}|g(\xi)|^{2}\mathrm{d}\xi.$ (87)
Plugging this into our expression for the square coefficient sum, we obtain
$\sum_{\ell,n}\int_{[0,2^{-\ell}e^{|n|\delta}]^{2}}\left|\sum_{s}\widehat{f}(\xi+2^{-\ell}e^{|n|\delta}s)\overline{\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s)}\right|^{2}\mathrm{d}\xi.$ (88)
Expanding the square now gives
$\sum_{\ell,n}\int_{[0,2^{-\ell}e^{|n|\delta}]^{2}}\sum_{s,s^{\prime}}\widehat{f}(\xi+2^{-\ell}e^{|n|\delta}s)\overline{\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s)\widehat{f}(\xi+2^{-\ell}e^{|n|\delta}s^{\prime})}\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s^{\prime})\mathrm{d}\xi$ (89)
$\quad=\sum_{\ell,n,s}\int_{\mathbb{R}^{2}}\widehat{f}(\xi)\overline{\widehat{\psi}(2^{\ell}B_{n\delta}\xi)\widehat{f}(\xi+2^{-\ell}e^{|n|\delta}s)}\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s)\mathrm{d}\xi,$ (90)
where $s^{\prime}$ again ranges over $\mathbb{Z}^{2}$ and we have extended the domain of the integral in the last step to include the sum over $s$ and renamed $s^{\prime}\rightarrow s$.
Separating out the $s=(0,0)^{\text{T}}$ terms, we obtain
$\int_{\mathbb{R}^{2}}|\widehat{f}(\xi)|^{2}\mathrm{d}\xi\sum_{\ell,n}|\widehat{\psi}(2^{\ell}B_{n\delta}\xi)|^{2}$ (91)
$\quad+\sum_{\ell,n}\sum_{s\neq(0,0)^{\text{T}}}\int_{\mathbb{R}^{2}}\widehat{f}(\xi)\overline{\widehat{f}(\xi+2^{-\ell}e^{|n|\delta}s)\widehat{\psi}(2^{\ell}B_{n\delta}\xi)}\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s)\mathrm{d}\xi.$ (92)
We note that the eigenvalues of $B_{n\delta}$ are given by $e^{n\delta}$ and $e^{-n\delta}$, which means that $\|e^{|n|\delta}B_{n\delta}s\|\geq\|s\|\geq 1$. Since we require that the mother boostlet $\psi$ has a Fourier support contained in a disk of diameter one, the product $\overline{\widehat{\psi}(2^{\ell}B_{n\delta}\xi)}\widehat{\psi}(2^{\ell}B_{n\delta}\xi+e^{|n|\delta}B_{n\delta}s)$ must always be zero whenever $s\neq(0,0)^{\text{T}}$. The second sum is therefore zero and we are left with the identity
$\sum_{\ell,n,m}|\langle f,\psi_{\ell,n,m}\rangle|^{2}=\int_{\mathbb{R}^{2}}|\widehat{f}(\xi)|^{2}\mathrm{d}\xi\sum_{\ell,n}|\widehat{\psi}(2^{\ell}B_{n\delta}\xi)|^{2}.$ (93)
The same result applies if we replace $\psi$ with $\psi^{\ast}$, and combining them gives
$\sum_{\ell,n,m}|\langle f,\psi_{\ell,n,m}\rangle|^{2}+|\langle f,\psi_{\ell,n,m}^{\ast}\rangle|^{2}=\int_{\mathbb{R}^{2}}|\widehat{f}(\xi)|^{2}\mathrm{d}\xi\sum_{\ell,n}|\widehat{\psi}(2^{\ell}B_{n\delta}\xi)|^{2}+|\widehat{\psi}^{\ast}(2^{\ell}B_{n\delta}\xi)|^{2}.$ (94)
Substituting in Eq. (39) then gives
$A\|f\|^{2}\leq\sum_{\ell,n,m}|\langle f,\psi_{\ell,n,m}\rangle|^{2}+|\langle f,\psi_{\ell,n,m}^{\ast}\rangle|^{2}\leq B\|f\|^{2},$ (95)
as desired.